You want to buy a house worth $3,230,000. You choose to take out a 30-year fixed rate conforming mortgage, put a downpayment of $1.4 million and borrow the rest. The APR of the mortgage is 4.563% compounded monthly.

What is the Present Value of the overall amount of interest you have to pay?
Answer #1

Amount financed = 3,230,000- 1,400,000 = 1830000

monthly rate = 4.563 /12 = .38025%

number of months = 30*12=360

Monthly payment = Amount financed /PVA .38025%,360

              = 1830000/ 195.91113

                = 9340.97 per month

Interest = Total payment made -amount financed

             = [9340.97*360 ] -1830000

              = 3362749-1830000

             = 1532749